# Late-time Evolution and Instabilities of Tidal Disruption Disks

### v1

@@ Abstract
Observations of tidal disruption events (TDEs) on a timescale of years after the main flare show evidence of continued activity in the form of optical/UV emission, quasi-periodic eruptions, and delayed radio flares. Motivated by this, we explore the time evolution of these disks using semi-analytic models to follow the changing disk properties and feeding rate to the central black hole (BH). We find that thermal instabilities typically begin $\sim 150-250\,{\rm days}$ after the TDE, causing the disk to cycle between high and low accretion states for up to $\sim 10-20\,{\rm yrs}$. The high state is super-Eddington, which may be associated with outflows that eject $\sim 10^{-3}-10^{-1}\,M_{\odot}$ with a range of velocities of $\sim 0.03-0.3c$ over a span of a couple of days and produce radio flares. In the low state, the accretion rate slowly grows over many months to years as continued fallback accretion builds the mass of the disk. In this phase, the disk may reach luminosities of $\sim 10^{41}-10^{42}\,{\rm erg\,s^{-1}}$ in the UV as seen in some late-time observations. We highlight the importance of the iron-opacity “bump” at $\approx 2\times 10^{5}\,{\rm K}$ in generating sufficiently high luminosities. This work suggests that joint optical/UV observations with radio monitoring could be key for following the disk state as the radio flares are produced.
@@ 1 Introduction
The TDE is seen as a flare that lasts for many weeks to month, and can be observed across a wide range of electromagnetic wavelengths (often for the same event). In recent years it has become clear that TDE BHs can remain active for many years following the initial flare. One way this is seen is from the optical/UV emission that can persist from the TDE location (van Velzen et al., 2019; Mummery & Balbus, 2020; Nicholl et al., 2024). Another is the presence of X-ray flares called quasi-periodic eruptions (QPEs, Miniutti et al., 2019; Giustini et al., 2020; Chakraborty et al., 2021; Arcodia et al., 2021, 2022; Miniutti et al., 2023b; Arcodia et al., 2024). It was expected that there was a connection between TDEs and QPEs due to host galaxy similarities (Wevers et al., 2022), theoretical arguments (Linial & Metzger, 2023; Franchini et al., 2023), and persistent X-ray emission (Chakraborty et al., 2021; Miniutti et al., 2023a; Quintin et al., 2023), and this was confirmed by an optical TDE that years later produced QPEs (Nicholl et al., 2024).
This emission observed years after TDEs is generally thought to require a long lasting disk. This would naturally produce the persistent optical/UV mission, and in the case of QPEs, a disk would be needed for either the star-disk collision models (e.g., Linial & Metzger, 2023; Franchini et al., 2023) or in the disk instability models (e.g., Pan et al., 2022; Kaur et al., 2023). Shen & Matzner (2014) considered the viscous evolution of TDE accretion disks on a timescale of up to $\sim 10^{4}\,{\rm yrs}$. They generally found that after a few months to about a year, a thermal instability would transition the disk to a gas-pressure-dominated low accretion state. This might naturally explain the jet shutoff at $\sim 500\,{\rm days}$ from the jetted TDE candidate Swift J1644+57 (Levan et al., 2011; Zauderer et al., 2013), but it makes the late time optical/UV emission difficult to reconcile with the low expected accretion rate at these times. Nevertheless, groups have fit the late time optical/UV and constructed disk models for the QPEs either by treating the disk scaleheight as a free parameter (e.g. Mummery et al., 2024) or by using an alternative viscosity prescription (e.g., van Velzen et al., 2019, which we discuss in more detail below). Linial & Metzger (2024) also considered the evolution of TDE disks over long time scales, but focused on potential feeding from the ablation of the QPE-producing star.
@@
Motivated by these issues, we conduct a semi-analytic exploration of TDE accretion disks. In Section 2, we summarize our one-zone model, and in Section 3, we compare our calculations to previous work. This is done to confirm our numerical methods and also to highlight the changes to the evolution we find due to the different physics we employ. In Section 4, we summarize the variety of disk evolutions we expect as a function of different TDE parameters. We compare our results to late time UV observations and radio flare properties in Section 5, and we conclude in Section 6 with a summary of our results and a discussion of future work.
@@ 2 One-Zone Disk Model
To understand the evolution of an accretion disk left over from a TDE, we use a one-zone disk model (e.g., Metzger et al., 2008; Shen & Matzner, 2014). Such a disk is characterized by a total mass $M_{d}$, characteristic radius $R_{d}$ (which is roughly the outer radius of the disk), and disk angular momentum
@@
where $M_{\rm BH}$ is the BH mass. Such a one-zone model can be a good approximation since the majority of the disk mass and angular momentum are at $R_{d}$, the location which satisfies $t_{\nu}\sim t$, where $t_{\nu}$ is the viscous timescale. Since $t_{\nu}\sim r^{3/2}$, interior to $R_{d}$ the disk can be assumed to be in steady state. We next explain the basic ingredients included to solve for the time-dependent disk properties.
@@ 2.2 Time Evolution Equations
The disk is fed via fallback of material from TDE at a rate $\dot{M}_{\rm fb}$ as described above, which then accretes through the disk at a rate $\dot{M}$. Some fraction of this mass transport may go into an outflow (such as when the disk reaches super-Eddington rates), while the remaining mass accretes all the way down to the BH. Either way, the total mass loss of the disk at any time adds up to $\dot{M}$, and thus the differential equation
@@
describes the mass evolution.
The angular momentum evolution requires a little more careful consideration. Fallback adds angular momentum to the disk with a specific value of $j_{\rm fb}$, but it can also be removed via viscously driven outflows. These two processes are represented in the following differential equation that governs disk angular momentum,
$\frac{dJ_{d}}{dt}=j_{\rm fb}\dot{M}_{\rm fb}-C(GM_{\rm BH}R_{d})^{1/2}\dot{M},$ (9)
where the constant $C$ is determined by the torque exerted by the wind on the disk. Assuming that the outflow produces no net torque (e.g., Stone & Pringle, 2001), then the angular momentum losses are due to the specific angular momentum at each disk radius and $C=2/3$ (Kumar et al., 2008). Although in principle this factor can be higher if large scale magnetic fields help transport additional angular momentum loss, we simply use $C=2/3$ when $\dot{M}$ is super-Eddington and $C=0$ otherwise. There is additional angular momentum loss due to accretion by the BH, but this can be ignored since the inner radius of the disk is $\ll R_{d}$.
@@ 2.3 Disk Structure
The surface density is $\Sigma=\rho H$, and mass conservation through the disk leads to $\dot{M}_{\rm acc}=3\pi\nu\Sigma$, where $\nu$ is the viscosity. We parameterize the viscosity using the usual $\alpha$-disk model (Shakura & Sunyaev, 1973)
@@
We ignore energy lost to the wind, although see the Appendix of Shen & Matzner (2014) for a discussion of the small correction from this effect.
@@
Figure 3: The black dashed line represents the surface density $\Sigma$ and temperature $T$ for equilibrium disk solutions using $M_{\rm BH}=10^{6}\,M_{\odot}$, $\alpha=0.1$, a fixed radius of $R_{d}=3\times 10^{13}\,{\rm cm}$, and varying $\dot{M}$ from high to low values going from top to bottom. Cooling beats heating on the left of these solutions, and heating beats cooling on the right. The red solid line shows an example disk evolution (the same as used in Figure 2). The disk starts from the bottom-left corner and evolve toward the right. Once it reaches the instability region, the disk cycles counter clockwise in this space, alternating between low and high states. In purple, we highlight where the iron-opacity bump impacts the evolution.
@@ 2.4 Fallback Heating
A potentially important physical effect that we do not include in our main models is the interaction of the fallback accretion with the spreading accretion disk. This collision can add additional heating to the disk and modify the energy balance given by Equation 13. How well the fallback stream is thermalized depend on the ram pressure of the incoming material and the pressure in the disk, so we save a detailed treatment of this for future work. Nevertheless, the heating rate should be proportional to $\dot{M}_{\rm fb}v_{\rm fb}^{2}/2$, where $v_{\rm fb}\approx(GM_{\rm BH}/R_{d})^{1/2}$. We use $R_{d}$ rather than $R_{c}$ for setting this velocity because the incoming stream is likely in a similar plane to the disk and thus will collide somewhere near the outer radius of the disk. Using this, we estimate the impact of including such effects by rewriting the energy equation as
@@
In Figure 4, we recalculate the equilibrium solutions from Figure 3, but now including fallback heating using the updated energy Equation (15) with $\eta=1$ and varying the value of $\dot{M}_{\rm fb}$. This shows that when $\dot{M}_{\rm fb}=M_{\odot}\,{\rm yr^{-1}}$, for this particular example, the instability no longer occurs. As we decrease $\dot{M}_{\rm fb}$, the instability reappears, and at $\dot{M}_{\rm fb}=10^{-6}\,M_{\odot}\,{\rm yr^{-1}}$, the equilibrium solutions are basically the same as not including fallback heating at all (the black dashed line). Note that for this treatment $\eta$ and $\dot{M}_{\rm fb}$ are degenerate with one another, so the exact fallback value where the instability disappears depends on $\eta$. We conducted a series of time evolution calculations with different levels of heating (not presented in this work) and found that for high $\eta$ the instability can be removed at early times, which causes $\dot{M}\approx\dot{M}_{\rm fb}$. Since the early evolution depends sensitively on the choice of $\eta$, we save a more detailed investigation of fallback heating for future work and focus on the late time evolution for our study here.
@@ 3 Comparisons to Previous Work
To find the disk evolution, we solve the differential Equations (8) and (9) explicitly, where $\dot{M}$ is set using energy balance as given by Equation (13) and described in Section 2.3. In Appendix A, we provide more details about how we solve for the disk structure and time evolution. To check our approach, in this section we make comparisons in some simplified limits.
To facilitate comparisons to Shen & Matzner (2014), we had to update some of the prescriptions we use in our disk model (our work is set to be consistent with Frank et al., 2002). To quickly summarize the differences, these are $\Sigma=2\rho H$, $\nu=2\alpha c_{s}H/3$, $Q^{+}=9\nu\Sigma\Omega^{2}/4$, $Q^{-}_{\rm rad}=4acT^{4}/(3\kappa\Sigma)$, and setting $\xi=1$. We use these values and relations for the remainder of this section, but for all other calculations shown in this work we use the prescriptions summarized in Section 2.
@@
In addition, Shen & Matzner (2014) use the analytic fallback rate give in Equation (3) rather than the numerical fallback rates we show in Figure 1. Integrating Equation (3) from $t=t_{\rm fb}$ to $t=\infty$ results in a mass of $103M_{*}/$ 10. They assume that an additional mass of $M_{*}/5$ circularizes during the time $\sim t_{\rm fb}$, so that basically $\sim t_{\rm fb}\times\dot{M}_{\rm fb}(t_{\rm fb})$ sets the initial mass of the disk. This then sums to give a total fallback mass of $M_{*}/2$ as is well known.
@@
where $L_{\rm Edd}$ is the Eddington luminosity, to mimic their Figure 7 as closely as possible. Their “fiducial model” (blue solid line) uses $M_{\rm BH}=10^{6}\,M_{\odot}$, $M_{*}=M_{\odot}$, $\beta=1$, $\alpha=0.01$, and purely $\kappa=\kappa_{\rm es}$ for the opacity (see the upper left panel in their Figure 7). We get an $R_{d}\propto t^{2/3}$ at early times like Shen & Matzner (2014). The accretion rate and radius then drop dramatically at $t\approx 10^{-0.6}\,{\rm yrs}$ due to the thermal instability. The accretion rate slowly increases as the disk builds from fallback accretion. At late times once $\dot{M}_{\rm fb}\ll\dot{M}$, the disk obeys the power-law behavior $\dot{M}\propto t^{-19/16}$ and $R_{d}\propto t^{3/8}$ as derived by Shen & Matzner (2014).
Figure 7: Evolution of the accretion rate $\dot{M}$, disk mass $M_{d}$, and disk radius $R_{d}$ over the first year using $M_{*}=M_{\odot}$, $\beta=1.85$, and $\alpha=0.1$, for three different values of $M_{\rm BH}$ as indicated. The black lines delineate the fallback accretion rate $\dot{M}_{\rm fb}$.
The other models in Figure 5 explore what happens as we change the model parameters away from the fiducial values used by Shen & Matzner (2014). First, we use OPAL opacities rather than strictly electron scattering (red dashed lines). The evolution is mostly the same, with the main differences being the late-time radius and an increased $\dot{M}$ during the low state. Next, we increase the viscosity to $\alpha=0.1$ and use the analytic opacity from Equation (14) (green dot-dashed lines). The higher viscosity makes the disk evolve more quickly, increasing the disk radius at early times and then causing the disk to go through multiple cycles of low and then high states for a few years until $\dot{M}_{\rm fb}\lesssim\dot{M}$. We note that such cycles were also seen in a subset of models explored by Shen & Matzner (2014), although this was not the focus on their work. Finally, we use $\alpha=0.1$ with the full OPAL opacities (purple dotted line). Although the high states now appear at different times in comparison to the previous model, the evolution is qualitatively similar. An interesting detail is that with the OPAL opacities $\dot{M}$ is slightly larger as is rises toward a high state, which will prove important when discussing late-time UV emission from TDE disks.

### v2

@@ Abstract
Observations of tidal disruption events (TDEs) on a timescale of years after the main flare show evidence of continued activity in the form of optical/UV emission, quasi-periodic eruptions, and delayed radio flares. Motivated by this, we explore the time evolution of these disks using semi-analytic models to follow the changing disk properties and feeding rate to the central black hole (BH). We find that thermal instabilities typically begin $\sim 100\,{\rm days}$ after the TDE, causing the disk to cycle between high and low accretion states for up to $\sim 10\,{\rm yrs}$. The high state is super-Eddington, which may be associated with outflows that eject $\sim 10^{-3}-10^{-1}\,M_{\odot}$ over $\sim 1-2\,{\rm days}$ with a range of velocities of $\sim 0.03-0.3c$. Collision between these mass ejections may cause radio flares. In the low state, the accretion rate slowly grows over months to years as continued fallback accretion builds the disk’s mass. In this phase, the disk has a luminosity of $\sim 10^{41}-10^{42}\,{\rm erg\,s^{-1}}$ in the optical/UV as seen in some late-time observations. Although the accretion cycles we find occur for a typical $\alpha$-disk, in nature the disk could be stabilized by other effects such as the disk’s magnetic field or heating from fallback accretion, the latter of which we explore. Thus higher cadence optical/UV observations along with joint radio monitoring will be key for following the disk state and testing these models.
@@ 1 Introduction
The TDE is seen as a flare that lasts for many weeks to months, and can be observed across a wide range of electromagnetic wavelengths (often for the same event). In recent years it has become clear that TDE BHs can remain active for many years following the initial flare. One way this is seen is from the optical/UV emission that can persist from the TDE location (van Velzen et al., 2019; Mummery & Balbus, 2020; Nicholl et al., 2024). Another is the presence of X-ray flares called quasi-periodic eruptions (QPEs, Miniutti et al., 2019; Giustini et al., 2020; Chakraborty et al., 2021; Arcodia et al., 2021, 2022; Miniutti et al., 2023b; Arcodia et al., 2024). It was expected that there was a connection between TDEs and QPEs due to host galaxy similarities (Wevers et al., 2022), theoretical arguments (Linial & Metzger, 2023; Franchini et al., 2023), and persistent X-ray emission (Chakraborty et al., 2021; Miniutti et al., 2023a; Quintin et al., 2023), and this was confirmed by an optical TDE that years later produced QPEs (Nicholl et al., 2024).
This emission observed years after TDEs is generally thought to require a long lasting disk. This would naturally produce the persistent optical/UV mission, and in the case of QPEs, a disk would be needed for either the star-disk collision models (e.g., Linial & Metzger, 2023; Franchini et al., 2023) or in the disk instability models (e.g., Pan et al., 2022; Kaur et al., 2023). Shen & Matzner (2014) considered the viscous evolution of TDE accretion disks on a timescale of up to $\sim 10^{4}\,{\rm yrs}$. They generally found that after a few months to about a year, a thermal instability would transition the disk to a gas-pressure-dominated low accretion state. This might naturally explain the jet shutoff at $\sim 500\,{\rm days}$ from the jetted TDE candidate Swift J1644+57 (Levan et al., 2011; Zauderer et al., 2013), but it makes the late time optical/UV emission difficult to reconcile with the low expected accretion rate at these times. Nevertheless, groups have fit the late time optical/UV and constructed disk models for the QPEs either by treating the disk scaleheight as a free parameter (e.g., Mummery et al., 2024) or by using an alternative viscosity prescription (e.g., van Velzen et al., 2019, which we discuss in more detail below). Lu (2022) updated the models from Shen & Matzner (2014) by implementing a more detailed opacity treatment and a higher viscosity. Linial & Metzger (2024) also considered the evolution of TDE disks over long time scales, but focused on potential feeding from the ablation of the QPE-producing star.
@@
The r- and g-band light curves years after TDEs show variations as well (Mummery et al., 2024). These are higher cadence than the UV data, which makes them more sensitive to detecting these changes, and generally show variations by a factor of $\sim 3$. Some events exhibit even larger variations, for example AT2021mhg and AT2020riz show an increase by a factor of $\sim 10$, while AT2018lni shows variations in the g-band by a factor of $\sim 6$. This all points to the late time emission not being as constant as may have been previously believed.
Motivated by these issues, we conduct a semi-analytic exploration of TDE accretion disks. In Section 2, we summarize the one-zone model we use in this work. This includes spelling out the guiding equations that are used to solve for the time evolution of the disk. We also explore the impact of a more detailed treatment of the opacity, comparing pure electron scattering to Kramers’ and OPAL opacities (Iglesias & Rogers, 1996). We show that our disk models cycle between low and high accretion states and discuss the physics that determines this and whether additional heating sources may stifle this instability. In Section 3, we compare our calculations to the work of Shen & Matzner (2014). This is done to confirm the numerical methods we employ (which are discussed in more detail in Appendix A), and also to highlight the changes to the evolution we find due to differences in the treatment of the physics. We show that the higher opacities and viscosities we use are key for finding the accretion cycles that were not highlighted before (similar to the conclusions of Lu, 2022).
We continue in Section 4 by summarizing the variety of disk evolutions we expect as a function of different TDE parameters. This shows that the accretion cycles happen more quickly for lower mass BHs. The high accretion states are super-Eddington and may eject $\sim 10^{-3}-10^{-1}\,M_{\odot}$ on a timescale of $\sim 1-2\,{\rm days}$. We compare our results to late time observations Section 5, showing that in the low accretion state our models can match the $\sim 10^{42}\,{\rm erg\,s^{-1}}$ luminosities seen in the UV and $\sim 10^{41}\,{\rm erg\,s^{-1}}$ luminosities seen in the optical. Furthermore, we expect smaller variations in the optical bands than in higher energy bands, and the factor of $\sim 3$ variations we find roughly match observations. We conclude in Section 6 with a summary of our results and a discussion of future work.
@@ 2 One-Zone Disk Model
To understand the evolution of an accretion disk left over from a TDE, we use a one-zone disk model (e.g., Metzger et al., 2008; Shen & Matzner, 2014; Lu, 2022). Such a disk is characterized by a total mass $M_{d}$, characteristic radius $R_{d}$ (which is roughly the outer radius of the disk), and disk angular momentum
@@
where $M_{\rm BH}$ is the BH mass. Such a one-zone model can be a good approximation since the majority of the disk mass and angular momentum are at $R_{d}$, the location which satisfies $t_{\nu}\sim t$, where $t_{\nu}$ is the viscous timescale. Since $t_{\nu}\sim r^{3/2}$, interior to $R_{d}$ the disk can be assumed to be in steady state. In Section 6, we come back to this assumption and discuss whether a one-dimensional model may be needed in some cases. We next explain the basic ingredients included to solve for the time-dependent disk properties.
@@ 2.2 Time Evolution Equations
The disk is fed via fallback of material from TDE at a rate $\dot{M}_{\rm fb}$ as described above, which then accretes through the disk at a rate $\dot{M}$. Some fraction of this mass transport may go into an outflow, while the remaining mass accretes all the way down to the BH. The outflow is expected to be launched within the spherization radius $R_{\rm sph}$, where the local accretion luminosity $GM_{\rm BH}\dot{M}/R_{\rm sph}$ exceeds the the Eddington luminosity $4\pi GM_{\rm BH}c/\kappa$ (Begelman, 1979),
$R_{\rm sph}=\min\left(R_{d},\frac{\kappa\dot{M}}{4\pi c}\right),$ (8)
where we use an electron scattering opacity for $\kappa$. Following previous studies (e.g., Metzger et al., 2008; Yuan & Narayan, 2014; Hu et al., 2022), we assume these disk outflows cause the mass inflow rate to decrease within $R_{\rm sph}$ as it approaches the BH (e.g., Blandford & Begelman, 1999)
$\dot{M}(r)=\left(\frac{r}{R_{\rm sph}}\right)^{p}\dot{M},$ (9)
where $10<p<$ 1 is a parameter that controls the strength of the outflow. The total mass outflow rate is thus
$\dot{M}_{\rm out}=\left[1-\left(\frac{R_{i}}{R_{\rm sph}}\right)^{p}\right]\dot{M},$ (10)
where $R_{i}$ is the inner radius of the disk. Even with an outflow, the total mass loss of the disk at any time adds up to $\dot{M}$, and thus the differential equation
@@
describes the mass evolution. This expression assumes that the mass feeding rate to the disk tracks the fallback rate. If stream collisions are responsible for circularizing material into a disk to begin with (as opposed to, e.g. nozzle shocks, see Steinberg & Stone, 2024), Lense-Thirring precession by a rapidly spinning BH can delay this process by causing the stream to miss colliding with itself for many orbits (Dai et al., 2013; Guillochon & Ramirez-Ruiz, 2015; Jankovifç et al., 2024). The details of this depend on the general relativistic precession versus the hydrodynamic evolution of the stream thickness. The circularization process still need further exploration, so by simply using $\dot{M}_{\rm fb}$ our results are focused on the period after material has begun to circularize into a disk.
We next consider the disk’s angular momentum evolution. Fallback adds angular momentum with a specific value of $j_{\rm fb}$, but it can also be removed via the outflows. These two processes are represented in the following differential equation,
$\frac{dJ_{d}}{dt}=j_{\rm fb}\dot{M}_{\rm fb}-C(GM_{\rm BH}R_{\rm sph})^{1/2}\dot{M}_{\rm out},$ (12)
where the constant $C$ is determined by the torque exerted by the wind on the disk. Assuming that the outflow produces no net torque (e.g., Stone & Pringle, 2001), then the angular momentum losses are due to the specific angular momentum at each disk radius, resulting in
$C=\frac{2p}{2p+1}$ (13)
(Kumar et al., 2008). In principle this factor can be higher if large scale magnetic fields help transport additional angular momentum loss, but we do not consider these effects in this work.
@@ 2.3 Disk Structure
The surface density is $\Sigma=\rho H$, and mass conservation through the disk leads to $\dot{M}=3\pi\nu\Sigma$, where $\nu$ is the viscosity. We parameterize the viscosity using the usual $\alpha$-disk model (Shakura & Sunyaev, 1973)
@@
We ignore energy lost to the wind, although see the Appendix of Shen & Matzner (2014) for a discussion of the small correction from this effect. We also do not include energy loss from vertical convection, which may be particularly important when the disk is super-Eddington (Jiang et al., 2013).
@@
Figure 3: The black dashed line represents the surface density $\Sigma$ and temperature $T$ for equilibrium disk solutions using $M_{\rm BH}=10^{6}\,M_{\odot}$, $\alpha=0.1$, a fixed radius of $R_{d}=3\times 10^{13}\,{\rm cm}$, and varying $\dot{M}$ from high to low values going from top to bottom. Cooling beats heating on the left of these solutions, and heating beats cooling on the right. The red solid line shows an example disk evolution. The disk starts from the bottom-left corner and evolve toward the right. Once it reaches the instability region, the disk cycles counter clockwise in this space, alternating between low and high states. In purple, we highlight where the iron-opacity bump impacts the evolution.
@@ 2.4 Fallback Heating
A potentially important physical effect that we do not include in our main models is the interaction of the fallback accretion with the spreading accretion disk. This collision can add additional heating to the disk and modify the energy balance given by Equation (17). How well the fallback stream is thermalized depends on the ram pressure of the incoming material, the pressure in the disk, and the trajectory of the fallback which is related to the hydrodynamics and dynamics of the stream and disk, so we save a detailed treatment of this for future work. Nevertheless, the heating rate should be proportional to $\dot{M}_{\rm fb}v_{\rm fb}^{2}/2$, where $v_{\rm fb}\approx(GM_{\rm BH}/R_{d})^{1/2}$. Note that a similar prescription for this heating was used by Lu (2022). We use $R_{d}$ rather than $R_{c}$ for setting this velocity, which assumes that the incoming stream is in a similar plane to the disk and thus will collide somewhere near the outer radius of the disk. Using this, we estimate the impact of including such effects by rewriting the energy equation as
@@
In Figure 4, we recalculate the equilibrium solutions from Figure 3, but now including fallback heating using the updated energy Equation (19) with $\eta=1$ and varying the value of $\dot{M}_{\rm fb}$. This shows that when $\dot{M}_{\rm fb}=M_{\odot}\,{\rm yr^{-1}}$, for this particular example, the instability no longer occurs. As we decrease $\dot{M}_{\rm fb}$, the instability reappears, and at $\dot{M}_{\rm fb}=10^{-6}\,M_{\odot}\,{\rm yr^{-1}}$, the equilibrium solutions are basically the same as not including fallback heating at all (the black dashed line).
Note that for this treatment $\eta$ and $\dot{M}_{\rm fb}$ are degenerate with one another, so the exact fallback value where the instability disappears depends on $\eta$. We conducted a series of time evolution calculations with different levels of heating (not presented in this work) and found that for high $\eta$ the instability can be removed at early times, which causes $\dot{M}\approx\dot{M}_{\rm fb}$. This is similar to the results of Linial & Metzger (2024), who also found that the disk can be stabilized with an additional heating source, but in their case this was from the shock by the QPE-generating star colliding with the disk.
Bonnerot et al. (2021) studied stream interactions and found that most of the orbital energy is dissipated by shocks that are away from the disk if there is a strong stream self-crossing shock that diverts the fallback in a somewhat spherical manner. This suggests that $\eta\ll 1$, which would make disk stabilization from fallback difficult. This is not a solved problem though, and these uncertainties show that future work is needed to better understand the interactions between the fallback gas and an existing disk so that the amount of heating can be better assessed.
@@ 3 Comparisons to Previous Work
To find the disk evolution, we solve the differential Equations (11) and (12) explicitly, where $\dot{M}$ is set using energy balance as given by Equation (17) and described in Section 2.3. In Appendix A, we describe the numerical scheme used to solve for the disk structure and time evolution. To check our approach, in this section we make comparisons in some simplified limits.
To facilitate comparisons to Shen & Matzner (2014), we had to update some of the prescriptions we use in our disk model (our work is set to be consistent with Frank et al., 2002). To quickly summarize the differences, these are $\Sigma=2\rho H$, $\nu=2\alpha c_{s}H/3$, $Q^{+}=9\nu\Sigma\Omega^{2}/4$, $Q^{-}_{\rm rad}=4acT^{4}/(3\kappa\Sigma)$, and setting $\xi=1$. We also assume here that all accretion above the Eddington rate is ejected rather than use the radially varying mass loss rate as described by Equation (10). We use these values and relations for the remainder of this section, but for all other calculations shown in this work we use the prescriptions summarized in Section 2.
@@
In addition, Shen & Matzner (2014) use the analytic fallback rate given in Equation (3) rather than the numerical fallback rates we show in Figure 1. Integrating Equation (3) from $t=t_{\rm fb}$ to $t=\infty$ results in a mass of $103M_{*}/$ 10. They assume that an additional mass of $M_{*}/5$ circularizes during the time $\sim t_{\rm fb}$, so that basically $\sim t_{\rm fb}\times\dot{M}_{\rm fb}(t_{\rm fb})$ sets the initial mass of the disk. This then sums to give a total fallback mass of $M_{*}/2$ as is well known.
@@
where $L_{\rm Edd}$ is the Eddington luminosity, to mimic their Figure 7 as closely as possible. Their “fiducial model” (blue solid line) uses $M_{\rm BH}=10^{6}\,M_{\odot}$, $M_{*}=M_{\odot}$, $\beta=1$, $\alpha=0.01$, and purely $\kappa=\kappa_{\rm es}$ for the opacity (see the upper left panel in their Figure 7). We get an $R_{d}\propto t^{2/3}$ at early times like Shen & Matzner (2014). The accretion rate and radius then drop dramatically at $t\approx 10^{-0.6}\,{\rm yrs}$ due to the thermal instability. The accretion rate slowly increases as the disk builds from fallback accretion. At late times once $\dot{M}_{\rm fb}\ll\dot{M}$, the disk obeys the power-law behavior $\dot{M}\propto t^{-19/16}$ and $R_{d}\propto t^{3/8}$ as derived by Cannizzo et al. (1990).
Figure 7: Evolution of the accretion rate $\dot{M}$, disk mass $M_{d}$, and disk radius $R_{d}$ over the first year using $M_{*}=M_{\odot}$, $\beta=1.85$, $\alpha=0.1$, and $p=0.5$ for three different values of $M_{\rm BH}$ as indicated. The black lines delineate the fallback accretion rate $\dot{M}_{\rm fb}$.
The other models in Figure 5 explore what happens as we change the model parameters away from the fiducial values used by Shen & Matzner (2014). First, we use OPAL opacities rather than strictly electron scattering (red dashed lines). The evolution is mostly the same, with the main differences being the late-time radius and an increased $\dot{M}$ during the low state. Next, we increase the viscosity to $\alpha=0.1$ and use the analytic opacity from Equation (18) (green dot-dashed lines). The higher viscosity makes the disk evolve more quickly, increasing the disk radius at early times and then causing the disk to go through multiple cycles of low and then high states for a few years until $\dot{M}_{\rm fb}\lesssim\dot{M}$. We note that such cycles were also seen in a subset of models explored by Shen & Matzner (2014), although this was not the focus on their work. Finally, we use $\alpha=0.1$ with the full OPAL opacities (purple dotted line). Although the high states now appear at different times in comparison to the previous model, the evolution is qualitatively similar. An interesting detail is that with the OPAL opacities $\dot{M}$ is slightly larger as it rises toward a high state.